# 22footnotetext: Corresponding author

### v1

@@ 3.2. Scanning
    - Three Directional Scan SegMamba (Xing et al., 2024)
    Pixelwise and Patchwise Scan LKM-UNet (Wang et al., 2024a)
    Omnidirectional Selective Scan VmambaIR (Shi et al., 2024)
    Hierarchical Scan Motion Mamba(Zhang et al., 2024b)
    Multi-Head Scan MHS-VM (Ji, 2024)
    Multi-Path Scan RSMamba(Chen et al., 2024a)
    3D BiDirectional Scan VideoMamba(Li et al., 2024a)
    Bidirectional Slice Scan (BSS) SliceMamba (Fan et al., 2024)
  7. (7)
    Efficient 2D Scan : Efficient Scan 2D (ES2D) (Pei et al., 2024) emphasizes efficient image scanning by skipping scan patches with a step size $p$. It partitions selected spatial dimension features into $m$ and $n$ using sine and cosine functions to determine the patch location. The entire operation is mathematically expressed in Equation 15.
    (15) $\textbf{{O}}_{i}\overset{scan}{\leftarrow}\textbf{{X}}[:,m::p,n::p],$
@@
    Multi-Path Scan : Multi-path scanning mechanism (Chen et al., 2024a) incorporates reverse, forward and random shuffling paths. A simple approach to combine the information flow from different paths would be averaging. However, the objective is to selectively activate the information from each path. Consequently, a gating mechanism is designed to manage the information flow from various paths.
  9. (9)
@@
    Omnidirectional Selective Scan : There are two information streams $\textbf{{F}}_{01},\textbf{{F}}_{02}\;\mathbf{\in}\;\mathbf{\mathbb{R}}^{B\times C\times H\times W}$, serving as inputs for the Omnidirectional Selective Scan (OSS) (Shi et al., 2024). In the first stream, bidirectional scanning is performed both longitudinally and transversely on $\textbf{{F}}_{01}$ to capture planar two-dimensional feature information. The second stream is refined by depthwise convolution and SiLU activation (Elfwing et al., 2017), capturing detailed patterns. These two streams are then fused within OSS, merging refined features with complementary information. After passing through a $1\times 1$ convolution, the output of OSS, $\textbf{{F}}_{OSS}\;\mathbf{\in}\;\mathbf{\mathbb{R}}^{B\times C\times H\times W}$ provides a detailed input representation, improving feature extraction and modeling capabilities.
  10. (10)
@@
    Hierarchical Scan : Hierarchical Temporal Mamba (HTM) (Zhang et al., 2024b) block processes compressed latent representations $z$ of dimensions (T, B, C) using a hierarchical scanning methodology. Initially, $z$ undergoes a linear projection to produce representations $x$ and $z$ of dimension $E$. A set of scans $K$ and memory matrices are applied, where each scan involves $1D$ convolution and linear projections to derive transformed outputs. These outputs are combined through a linear projection to produce the final transformed representations $\textbf{{z}}_{\textbf{{HTM}}}$, efficiently capturing diverse motion densities and minimizing computational overhead.
  11. (11)
@@
    Multi-Head Scan : Multi-headed scanning (Ji, 2024) processes an input embedding map $\textbf{{X}}_{l-1}$ with shape (B, H, W, C) to produce an output embedding map $\textbf{{X}}_{l}$ of the same shape through linear transformations and concatenations. Initially, the input is reshaped to (B, C, H, W ) and then processed with $n$ scan heads. Each scan head projects the input onto a subspace and routes it through $K$ scan routes, involving specific transformations, activation, and rearrangement. Scan route outputs are concatenated and fused using the coefficient of variation and summation, followed by ReLU. The combined results from all scan heads are concatenated along the channel axis and optionally projected back to the original number of channels. Finally, the output is reshaped to (B, H, W, C) and returned.
  12. (12)
@@
    Other Scanning Mechanisms: Segmamba (Xing et al., 2024) uses TOM, which stands for Tri-Orientated Spatial Mamba Block. This block employs scanning along three dimensions: height, width and channels and three directions: forward direction, reverse direction, and inter-slice direction. LKM-UNet (Wang et al., 2024a) applies a two-scan strategy where the first scan involves pixel-level scanning. Each pixel is scanned unidirectionally (forward) and max-pooled into a single image. Subsequently, another round of unidirectional scanning (forward) is performed on these pooled images. Bidirectional Slice Scan (BSS) (Fan et al., 2024) proposed in Slice Mamba plays a crucial role in its architecture. Firstly the spatial dimension height and width are split into $m$ and $n$ windows separately resulting in the sequence of shape $F^{h}=\{\mathrm{f}_{i}^{h}\in\mathbb{R}^{m\times W\times C}|i=1,2,...\frac{H}{m}\}$ horizontally and $F^{v}=\{\mathrm{f}_{j}^{v}\in\mathbb{R}^{H\times n\times C}|j=1,2,...\frac{W}{n}\}$ sequence of shape vertically. Scanning is applied both in horizontal (both forward and backward) and vertical ‘(both forward and backward) direction. The optimal size of $m$ and $n$ are chosen through adaptive slice search which uses Neural Architecture Search (NAS) for each layer of mamba blocks using a single path one shot.
3.3. Mamba Optimizations
@@
3.3.1. Lightweight and Efficient
Lightweight and efficient models are designed to be smaller, quicker, and use fewer resources, while maintaining good performance. Table 2 compares the above mentioned lightweight models based on Giga Floating-point Operations Per Second (GFLOPs), number of parameters, and FPS Frames Per Second (FPS). These metrics provide a detailed evaluation of each model’s computational efficiency, complexity and speed respectively.
Light Mamba UNet (LightM UNet) proposed by Liao et al. (Liao et al., 2024) combines Mamba and UNet architectures in a lightweight framework which aims to tackle computational challenges in real world medical environment. The Residual Vision Mamba (RVM) layer is proposed to improve SSM for deep semantic feature extraction from images in a pure Mamba-based manner.
LightM-UNet overcomes the existing state-of-the-art methods with only 1.09 M parameters and 267.19 GFLOPs. UltraLight Vision Mamba UNet (UltraLight VM-UNet) introduced by Wu et al. (Wu et al., 2024b) is a lightweight vision Mamba model. An excellent performance is achieved by Parallel vision Mamba (PVM) method that is used for efficiently processing deep features with lowest computational complexity, while maintaining the overall number of processing channels constant. PVM is primarily composed using Mamba combined with residual connections and adjustment factors. This combination allows traditional Mamba to capture remote spatial relations without introducing additional parameters and computational complexity. Several comparisons have been done with the state-of-the-art lightweight models across three public skin lesion datasets. In this analysis, the UltraLight VM-UNet shows the competitive performance with only 0.049M parameters and 0.060 GFLOPs. UltraLight VM-UNet parameters are 87.84% lower than the parameters of LightM-UNet.
Table 2. Comparison of Lightweight Models based on Parameters, GFLOPs, and FPS
@@ 3.3.1. Lightweight and Efficient
LightM UNet (Liao et al., 2024) 1.09 267.19 -
UltraLight VM-UNet (Wu et al., 2024b) 0.049 0.060 -
MUCM-Net (Yuan et al., 2024) 0.071 to 0.139 0.055 to 0.064 -
LightCF-Net (Ji et al., 2024) 1.52 3.25 33
MiM-ISTD (Chen et al., 2024c) 4.76 3.95 -
@@
MUCM-Net proposed by Yuan et al. (Yuan et al., 2024) is an efficient model which combines Mamba State-Space Models with UCM-Net architecture to improve segmentation and feature learning. In this model, Mamba-UCM is optimized for mobile deployment, providing high accuracy with minimal computational requirements (approximately $0.055-0.064GFLOPs$ and $0.071-0.139M$ parameters).
@@
LightCF-Net proposed by Ji et al. (Ji et al., 2024) is a novel and efficient lightweight architecture used as a long-range context fusion network for real-time polyp segmentation. A new FAEncoder module has been developed, merging Large Kernel Attention (LKA) with channel attention mechanisms to extract deep representational features of polyps and uncover long-range relationships. Furthermore, a novel Visual Attention Mamba Module (VAM) module has been integrated into skip connections to capture extensive contextual dependencies from encoder-extracted features, prioritizing crucial information and mitigating background noise interference using attention mechanisms. Evaluation is done on four polyp segmentation datasets which showcases its operational efficiency and segmentation accuracy compared to leading lightweight polyp segmentation networks. While this method achieves good results in segmentation task, the complexity of the medical environment and limitations of labeled data during training pose challenges, preventing it from fully meeting the demands of medical applications. The model features 1.52 M parameters, operates at 3.25 GFLOPs, and achieves a frame rate of 33 FPS. Chen et al. (Chen et al., 2024c) proposed MiM-ISTD for Infrared Small Target Detection (ISTD). It utilizes Mamba to effectively capture both local and global information from the given data. This approach ensures higher efficiency with very less computational costs. Experiments conducted on NUAA-ISTD and IRSTD-1k datasets, where MiM-ISTD demonstrated superior performance in terms of both accuracy and efficiency compared to other related methods. MiM-ISTD is ten times faster than the state-of-the-art method and it reduced GPU memory usage by 73.4% during high resolution image testing.
3.4. Techniques and Adaptations
@@
In this section, we explore techniques and adaptations for Mamba architectures such as weakly supervised, semi-supervised ands self-supervised approaches. These approaches are used in scenarios where data annotations are absent, partially present or inconsistent, and used to improve the model’s ability to learn from unstructured or incomplete or semi-structured data.
3.4.1. Weakly Supervised Learning
@@
Weakly Supervised Learning (WSL) uses a small amount of correctly labeled data along with a large amount of data with incomplete labels. Instead of having detailed labels for each data, this approach works with data that has noisy and partial labels. Wang et al. (Wang and Ma, 2024) proposes Weak-Mamba-UNet, a WSL strategy which incorporates three different architectures but with the same symmetrical encoder-decoder networks. The three networks consist of a CNN based U-Net, known for capturing local features; a Swin Transformer-based SwinUNet, which excels in understanding global context and a VMamba-based Mamba-UNet, for efficiently capturing long-range dependency. The proposed WSL framework employs a multi-view cross-supervised learning approach for scribble-based supervised medical image segmentation. The work introduced partial cross-entropy to leverage only the scribble annotations during the training of the network. The overall loss is composed of the scribble-based partial cross-entropy loss and the dense-signal pseudo label dice-coefficient loss. The network demonstrates promising results in segmentation tasks achieving an accuracy of 99.63% on MRI Cardiac (Bernard et al., 2018) dataset.
3.4.2. Semi-Supervised Learning
@@
Semi-supervised learning uses a fewer amount of labeled data and a larger amount of unlabeled data during training. Ma et al. (Ma and Wang, 2024) proposes Semi-Mamba-UNet, a semi-supervised learning framework integrated with mamba-based segmentation network. It supports the complementary strengths of Mamba-UNet and UNet which uses labeled and a large amount of unlabeled data respectively. A Pixel-Level Contrastive learning strategy is proposed to increase feature learning from a pair of projectors. A network via pseudo labeling is used to train other network using a pixel-level cross-supervised learning strategy. The overall loss is the sum of three components: supervision loss, self-supervised contrastive loss, and semi-supervised loss. Semi-Mamba-UNet was tested on ACDC MRI Cardiac Dataset (Bernard et al., 2018) using 5% and 10% labeled data, along with remaining unlabeled data. With 5% labeled data, it achieved a dice coefficient of 0.8386, accuracy rate of 0.9936, sensitivity of 0.7992, and specificity of 0.9483. Hausdorff Distance (HD) was 6.2139, and Average Surface Distance (ASD) was 1.6406. When trained with 10% labeled data, the model’s performance improved, achieving a dice coefficient of 0.9114, accuracy rate of 0.9964, sensitivity of 0.9146, specificity of 0.9821, HD of 3.9124, and ASD of 1.1698.
3.4.3. Self Supervised Learning
@@
Self-supervised learning is a method where the model learns from unlabeled data by creating its own labels. Instead of depending on external manually annotated labels, the model generates these labels internally based on data itself. Nasiri et al. (Nasiri-Sarvi et al., 2024) proposes Vim4Path which uses Vision Mamba within DINO by Caron et al. (Caron et al., 2021) for representation learning. The research aims to explore and adapt ViM to use in SSL. DINO, a well known self-supervised learning framework employs self-distillation in a teacher-student setup, where both networks share identical architectures but have different parameters. The study compares two architectures in both slide-level and patch-level classification tasks using Camelyon16 dataset for benchmarking purposes. CLAM framework by Lu et al. (Lu et al., 2020) is employed to compare various architectures for slide-level classification. It utilizes attention-based multiple-instance learning, enabling the identification of sub-regions within slides that are most indicative of the slide-level label. This approach allows the model to focus on the most relavent features without need for detailed annotation. Zhou et al. (Zhou et al., 2024) proposes MGI, a new multimodal model which uses both genetic and image data. A self-supervised contrastive learning strategy is used during pre-training to align visual encoder and gene encoder on paired genetic and image data, allowing visual encoder to learn relevant features from a genetic perspective. This process is followed by a lightweight multimodal attention fusion decoder that integrates image and gene data. Using mamba to extract features from gene data addresses the issues faced by previous encoders in capturing long-range dependencies in long gene sequences. Tang et al. (Tang et al., 2024b) proposed MambaMIM, a 3D-UNet-based architecture for self-supervised learning. MambaMIM integrates 3D sparse convolution with Mamba blocks in the encoder. It introduces Selective Structured State Space Sequence Token-interpolation (S6T), which generates interpolated vectors between consecutive input vectors, $y_{i}$ and $y_{i+1}$. The interpolated sequence is processed through a linear layer before the decoder block. In the decoder, the unmasked sequence from the encoder is also interpolated using S6T, while sparse features are filled with learnable tokens, transforming them into dense features for upsampling.
@@
Contrastive learning is a technique within self-supervised learning that focuses on learning representations by comparing pairs of data samples. Yang et al. (Yang et al., 2024b) proposes Contrastive Masked Vim autoencoder (CMViM) which is the efficient representation learning method for 3D multi-modal data. To reconstruct 3D masked multi-modal data, it incorporates ViM into mask encoder so that it can effectively capture long range dependencies in 3D medical data. To align multi-modal representations, it introduces intramodal and intermodal contrastive learning mechanisms. Intra-modal contrastive learning module is introduced to capture discriminative features within the same modality. Inter-modal contrastive learning mechanism is introduced to align cross-modality representations from different modalities. CMViM outperforms other state-of-the-art methods in Alzheimer’s disease diagnosis. Ma et al. (Ma and Wang, 2024) introduces a pixel level contrastive learning strategy for feature learning maximization from representations that are projected in a pair of projectors.
3.4.4. Multimodal Learning
Figure 19. Visualization of Fusion Mamba Framework (Xie et al., 2024a) and its Main Components

### v3

@@ 3.2 Scanning
    - Three Directional Scan SegMamba [16]
    Pixelwise and Patchwise Scan LKM-UNet [18]
    Omnidirectional Selective Scan VmambaIR [103]
    Hierarchical Scan Motion Mamba[104]
    Multi-Head Scan MHS-VM [105]
    Multi-Path Scan RSMamba[106]
    3D BiDirectional Scan VideoMamba[107]
    Bidirectional Slice Scan (BSS) SliceMamba [108]
  7. 7.
    Efficient 2D Scan : Efficient Scan 2D (ES2D) [101] emphasizes efficient image scanning by skipping scan patches with a step size $p$. It partitions selected spatial dimension features into $m$ and $n$ using sine and cosine functions to determine the patch location. The entire operation is mathematically expressed in Equation 15.
    $\textbf{{O}}_{i}\overset{scan}{\leftarrow}\textbf{{X}}[:,m::p,n::p],$ (15)
@@
    Multi-Path Scan : Multi-path scanning mechanism [106] incorporates reverse, forward and random shuffling paths. A simple approach to combine the information flow from different paths would be averaging. However, the objective is to selectively activate the information from each path. Consequently, a gating mechanism is designed to manage the information flow from various paths.
  9. 9.
    Omnidirectional Selective Scan : There are two information streams
@@
    $\textbf{{F}}_{01},\textbf{{F}}_{02}\;\mathbf{\in}\;\mathbf{\mathbb{R}}^{B\times C\times H\times W}$, serving as inputs for the Omnidirectional Selective Scan (OSS) [103]. In the first stream, bidirectional scanning is performed both longitudinally and transversely on $\textbf{{F}}_{01}$ to capture planar two-dimensional feature information. The second stream is refined by depthwise convolution and SiLU activation [146], capturing detailed patterns. These two streams are then fused within OSS, merging refined features with complementary information. After passing through a $1\times 1$ convolution, the output of OSS, $\textbf{{F}}_{OSS}\;\mathbf{\in}\;\mathbf{\mathbb{R}}^{B\times C\times H\times W}$ provides a detailed input representation, improving feature extraction and modeling capabilities.
  10. 10.
@@
    Hierarchical Scan : Hierarchical Temporal Mamba (HTM) [104] block processes compressed latent representations $z$ of dimensions (T, B, C) using a hierarchical scanning methodology. Initially, $z$ undergoes a linear projection to produce representations $x$ and $z$ of dimension $E$. A set of scans $K$ and memory matrices are applied, where each scan involves $1D$ convolution and linear projections to derive transformed outputs. These outputs are combined through a linear projection to produce the final transformed representations $\textbf{{z}}_{\textbf{{HTM}}}$, efficiently capturing diverse motion densities and minimizing computational overhead.
  11. 11.
@@
    Multi-Head Scan : Multi-headed scanning [105] processes an input embedding map $\textbf{{X}}_{l-1}$ with shape (B, H, W, C) to produce an output embedding map $\textbf{{X}}_{l}$ of the same shape through linear transformations and concatenations. Initially, the input is reshaped to (B, C, H, W ) and then processed with $n$ scan heads. Each scan head projects the input onto a subspace and routes it through $K$ scan routes, involving specific transformations, activation, and rearrangement. Scan route outputs are concatenated and fused using the coefficient of variation and summation, followed by ReLU. The combined results from all scan heads are concatenated along the channel axis and optionally projected back to the original number of channels. Finally, the output is reshaped to (B, H, W, C) and returned.
  12. 12.
@@
    Other Scanning Mechanisms: Segmamba [16] uses TOM, which stands for Tri-Orientated Spatial Mamba Block. This block employs scanning along three dimensions: height, width and channels and three directions: forward direction, reverse direction, and inter-slice direction. LKM-UNet [18] applies a two-scan strategy where the first scan involves pixel-level scanning. Each pixel is scanned unidirectionally (forward) and max-pooled into a single image. Subsequently, another round of unidirectional scanning (forward) is performed on these pooled images. Bidirectional Slice Scan (BSS) [108] proposed in Slice Mamba plays a crucial role in its architecture. Firstly the spatial dimension height and width are split into $m$ and $n$ windows separately resulting in the sequence of shape $F^{h}=\{\mathrm{f}_{i}^{h}\in\mathbb{R}^{m\times W\times C}|i=1,2,...\frac{H}{m}\}$ horizontally and $F^{v}=\{\mathrm{f}_{j}^{v}\in\mathbb{R}^{H\times n\times C}|j=1,2,...\frac{W}{n}\}$ sequence of shape vertically. Scanning is applied both in horizontal (both forward and backward) and vertical ‘(both forward and backward) direction. The optimal size of $m$ and $n$ are chosen through adaptive slice search which uses Neural Architecture Search (NAS) for each layer of mamba blocks using a single path one shot.
3.3 Mamba Optimizations
@@
3.3.1 Lightweight and Efficient
Lightweight and efficient models are designed to be smaller, quicker, and use fewer resources, while maintaining good performance. Light Mamba UNet (LightM UNet) proposed by Liao et al. [80] combines Mamba and UNet architectures in a lightweight framework which aims to tackle computational challenges in real world medical environment. The Residual Vision Mamba (RVM) layer is proposed to improve SSM for the deep semantic feature extraction from images in a pure Mamba-based manner. UltraLight Vision Mamba UNet (UltraLight VM-UNet) introduced by Wu et al. [81] is a lightweight vision Mamba model. An excellent performance is achieved by Parallel vision Mamba (PVM) method that is used for efficiently processing deep features with the lowest computational complexity, while maintaining the overall number of processing channels constant. PVM is primarily composed of Mamba combined with residual connections and adjustment factors. This combination allows traditional Mamba to capture remote spatial relations without introducing additional parameters and computational complexity.
Table 2: Comparison of Lightweight Models based on Parameters, GFLOPs, and FPS
@@ 3.3.1 Lightweight and Efficient
LightM UNet [80] 1.09 267.19 -
UltraLight VM-UNet [81] 0.049 0.060 -
MUCM-Net [90] 0.071 to 0.139 0.055 to 0.064 -
LightCF-Net [95] 1.52 3.25 33
MiM-ISTD [96] 4.76 3.95 -
@@
Table 2 compares the above mentioned lightweight models based on Giga Floating-point Operations Per Second (GFLOPs), number of parameters, and FPS Frames Per Second (FPS). These metrics provide a detailed evaluation of each model’s computational efficiency, complexity and speed respectively. MUCM-Net proposed by Yuan et al. [90] is an efficient model which combines Mamba State-Space Models with UCM-Net architecture to improve segmentation and feature learning. In this model, Mamba-UCM is optimized for mobile deployment, providing high accuracy with minimal computational requirements (approximately 0.0550.0640.055-0.064 GFLOPs and 0.0710.139M0.071-0.139M parameters). LightCF-Net proposed by Ji et al. [95] is a novel and efficient lightweight architecture used as a long-range context fusion network for real-time polyp segmentation. A new FAEncoder module has been developed, merging Large Kernel Attention (LKA) with channel attention mechanisms to extract deep representational features of polyps and uncover long-range relationships. Furthermore, a novel Visual Attention Mamba Module (VAM) module has been integrated into skip connections to capture extensive contextual dependencies from encoder-extracted features. Chen et al. [96] proposed MiM-ISTD for Infrared Small Target Detection (ISTD). It utilizes Mamba to effectively capture both local and global information from the given data. This approach ensures higher efficiency with very less computational costs.
From this Section 3.3.1, UltraLight VM-UNet emerges as the efficient model in terms of both size and speed, using just 0.049M parameters and 0.060 GFLOPs. It achieves this through its Parallel Vision Mamba (PVM) design, which handles deep feature processing effectively. By combining Mamba with residual connections, the model captures long-range spatial relationships without adding extra complexity. In comparison, the Transformer-based models [147, 148] are much larger and more demanding. This makes UltraLight VM-UNet a more practical choice for resource-constrained medical applications.
3.4 Techniques and Adaptations
@@
In this section, we explore techniques and adaptations for Mamba architectures such as weakly supervised, semi-supervised and self-supervised approaches. These approaches are used in scenarios where data annotations are absent, partially present or inconsistent, and used to improve the model’s ability to learn from unstructured or incomplete or semi-structured data.
3.4.1 Weakly Supervised Learning
@@
Weakly Supervised Learning (WSL) uses a small amount of correctly labeled data along with a large amount of data with incomplete labels. Instead of having detailed labels for each data, this approach works with data that have noisy and partial labels. Wang et al. [79] proposes Weak-Mamba-UNet, a WSL strategy which incorporates three different architectures but with the same symmetric encoder-decoder networks. The three networks consist of a CNN based U-Net, known for capturing local features; a Swin Transformer-based SwinUNet, which excels in understanding global context and a VMamba-based Mamba-UNet, for efficiently capturing long-range dependency. The proposed WSL framework employs a multi-view cross-supervised learning approach for scribble-based supervised medical image segmentation. The work introduced partial cross-entropy to leverage only the scribble annotations during the training of the network. The overall loss is composed of the scribble-based partial cross-entropy loss and the dense-signal pseudo label dice-coefficient loss.
3.4.2 Semi-Supervised Learning
@@
Semi-supervised learning uses a fewer amount of labeled data and a larger amount of unlabeled data during training. Ma et al. [84] proposes Semi-Mamba-UNet, a semi-supervised learning framework integrated with a mamba-based segmentation network. It supports the complementary strengths of Mamba-UNet and UNet which uses labeled and a large amount of unlabeled data respectively. A Pixel-Level Contrastive learning strategy is proposed to increase feature learning from a pair of projectors. A network via pseudo labeling is used to train other network using a pixel-level cross-supervised learning strategy. The overall loss is the sum of three components: supervision loss, self-supervised contrastive loss, and semi-supervised loss. Semi-Mamba-UNet was tested on ACDC MRI Cardiac Dataset [54] using 5% and 10% labeled data, along with the remaining unlabeled data.
3.4.3 Self Supervised Learning
@@
Self-supervised learning is a method in which the model learns from unlabeled data by creating its own labels. Instead of depending on external manually annotated labels, the model generates these labels internally based on the data itself. Nasiri et al. [22] proposes Vim4Path which uses Vision Mamba within DINO by Caron et al. [141] for representation learning. The research aims to explore and adapt ViM for use in SSL. DINO, a well known self-supervised learning framework employs self-distillation in a teacher-student setup, where both networks share identical architectures but have different parameters. The study compares two architectures in slide-level and patch-level classification tasks using the Camelyon16 dataset for benchmarking purposes. CLAM framework by Lu et al. [149] is employed to compare various architectures for slide-level classification. It utilizes attention-based multiple-instance learning, enabling the identification of sub-regions within slides that are most indicative of the slide-level label. This approach allows the model to focus on the most relavent features without the need for detailed annotation. Zhou et al. [91] propose MGI, a new multimodal model which uses genetic and image data. A self-supervised contrastive learning strategy is used during pre-training to align visual encoder and gene encoder on paired genetic and image data, allowing visual encoder to learn relevant features from a genetic perspective. Tang et al. [93] proposed MambaMIM, a 3D-UNet-based architecture for self-supervised learning. MambaMIM integrates 3D sparse convolution with Mamba blocks in the encoder. It introduces Selective Structured State Space Sequence Token-interpolation (S6T), which generates interpolated vectors between consecutive input vectors, $y_{i}$ and $y_{i+1}$. The interpolated sequence is processed through a linear layer before the decoder block. In the decoder, the unmasked sequence from the encoder is also interpolated using S6T, while sparse features are filled with learnable tokens, transforming them into dense features for upsampling.
@@
Contrastive learning is a technique within self-supervised learning that focuses on learning representations by comparing pairs of data samples. Yang et al. [94] proposes Contrastive Masked Vim autoencoder (CMViM) which is the efficient representation learning method for 3D multi-modal data. To reconstruct 3D masked multi-modal data, it incorporates ViM into mask encoder so that it can effectively capture long range dependencies in 3D medical data. To align multi-modal representations, it introduces intramodal and intermodal contrastive learning mechanisms. Intra-modal contrastive learning module is introduced to capture discriminative features within the same modality. Inter-modal contrastive learning mechanism is introduced to align cross-modality representations from different modalities. CMViM outperforms other state-of-the-art methods in Alzheimer’s disease diagnosis. Ma et al. [84] introduces a pixel level contrastive learning strategy for feature learning maximization from representations that are projected in a pair of projectors.
From this Section 3.4.3, we can see that the need for manual labeling is reduced by learning directly from unlabeled data. These models capture long-range dependencies in complex medical data such as images and gene sequences. Contrastive learning helps align features across different modalities, improving performance on classification and segmentation tasks without requiring detailed annotations. However, these models require careful design of architecture and large amounts of unlabeled data.
3.4.4 Multimodal Learning